OpenStudy (anonymous):

Factor 16j2 + 24j + 9.

OpenStudy (anonymous):

j2 is j to the second power

OpenStudy (anonymous):

yummydum u replying?

OpenStudy (anonymous):

Do you recognize that this is a perfect square trinomial? \[(a + b)^2 = a^2 + 2ab + b^2\]

OpenStudy (anonymous):

\[16j^2+24j+9\] multiply the first and last coefficient and find factors of the product that add up to the coefficient of the second term \[9\times 16=144\]\[12\times12=144\]\[12+12=24\] now split the middle term and write it using the two factors\[16j^2+12j+12j+9\]\[(16j^2+12j)+(12j+9)\]now find GCFs for the parenthesis \[4j(4j+3)+3(4j+3)\]^this can also be written as\[4j+3(4j+3)\] so the factored form of \[16j^2+24j+9\]is\[(4j+3)(4j+3)\]or\[(4j+3)^2\]

OpenStudy (anonymous):

@alittlehaphazard do you get it? :)

jimthompson5910 (jim_thompson5910):

Multiply 16 and 9 to get 144 Now list all the ways to multiply to 144 1*144 = 144 2*72 = 144 3*48 = 144 4*36 = 144 6*24 = 144 8*18 = 144 9*16 = 144 12*12 = 144 (-1)*(-144) = 144 (-2)*(-72) = 144 (-3)*(-48) = 144 (-4)*(-36) = 144 (-6)*(-24) = 144 (-8)*(-18) = 144 (-9)*(-16) = 144 (-12)*(-12) = 144 Of this somewhat large list, notice how 12 and 12 add to 24 AND multiply to 144 So we can break up 24j into 12j + 12j and factor by grouping 16j^2 + 24j + 9 16j^2 + 12j + 12j + 9 (16j^2 + 12j) + (12j + 9) 4j(4j+3) + (12j + 9) 4j(4j+3) + 3(4j + 3) (4j+3)(4j+3) (4j+3)^2 So 16j^2 + 24j + 9 completely factors to (4j+3)^2

OpenStudy (anonymous):

i wrote all that too and then realized that alittlehaphazard closed his question...

OpenStudy (anonymous):

haha thank u (:

OpenStudy (anonymous):

your welcome :)